PLEAS I NEED HELP
1.IF ONE OF THE ANGLES IN A RIGHT TRIANGLE MEASURES 45 DEGREES, WHAT ARE THE MEASURES OF THE OTHER TWO ANGLES?
2.FIND THE PERIMETER OF A TRIANGLE WITH SIDES OF 4M, 2.3 M AN 3.7 M
PLEASE HELP

Answer by sarah_adam(201)

in right triangle we know one angle is 90 degrees
but we also know that the sum of measures of the three angles in a triangle is 180 degrees
also given one angle is 45 degrees
so the measure of third angle is:
x + y + z = 180
let x = 90 and y = 45
90 + 45 + z = 180
z = 180 - 45 - 90 = 45
so the three angles are 45, 45, 90

Perimeter of triangle = a+ b + c
a = 4 m ;b = 2.3 m ;and c = 3.7m
perimeter = 4+2.3+3.7 = 10 m